# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance

### v1

@@ II-A Notations
The sets of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We identify vectors using bold lowercase letters. The Euclidean norm of a vector $\mathbf{p}\in\mathbb{R}^{n}$ is denoted by $\lVert\mathbf{p}\rVert$, and an Euclidean ball of radius $r\geq 0$ centered at $\mathbf{p}$ is represented by $\mathcal{B}_{r}(\mathbf{p})=\{\mathbf{q}\in\mathbb{R}^{n}|\lVert\mathbf{q}-\mathbf{p}\rVert\leq r\}.$ The set of $n-$dimensional unit vectors is given by $\mathbb{S}^{n-1}=\{\mathbf{p}\in\mathbb{R}^{n}|\lVert\mathbf{p}\rVert=1\}.$ The three-dimensional Special Orthogonal group is defined by $SO(3):=\{\mathbf{R}\in\mathbb{R}^{3\times 3}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{R}\mathbf{R}^{\top},\det(\mathbf{R})=1\}.$ For a given vector $\mathbf{x}:=[x_{1},x_{2},x_{3}]^{\top}\in\mathbb{R}^{3}$, we define $\mathbf{x}^{\times}$ as the skew-symmetric matrix, which is given by
@@ II-C Geometric subsets of $\mathbb{R}^{n}$
II-C1 Line segment
@@ IV-A Hybrid control design
$\underbrace{\begin{matrix}\mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}\;=\mathbf{0},\\ \;=0,\\ \;=0\\ \;=1,\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}}\quad\quad\underbrace{\begin{bmatrix}\mathbf{h^{+}}\\ \mathbf{a^{+}}\\ m^{+}\\ {s^{+}}\end{bmatrix}\begin{matrix}\;\in\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s),\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{J}}$ (11o)
where $\kappa_{s}>0$, $\kappa_{r}>0$, and $\xi=(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$. The variable $\mathbf{h}$ denotes the hit point, which is the location where the robot switches from the move-to-target mode to the obstacle-avoidance mode. The vector $\mathbf{a}\in\mathbb{S}^{2}$ is instrumental for the construction of the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$, used in (11a). The scalar variable $s\in\mathbb{R}_{\geq 0}$ allows the robot to switch from the obstacle-avoidance mode to the move-to-target mode when initially in the obstacle-avoidance mode. Details of this switching process are provided later in Section IV-B. The sets $\mathcal{F}$ and $\mathcal{J}$ are the flow and jump sets related to different modes of operation, respectively, whose constructions are provided in Section IV-B. The update law $\mathbf{L}$, which allows the robot to update the values of the variables $\mathbf{h}$, $\mathbf{a}$, and $m$ based on the current location of the robot with respect to the nearest obstacle and the target location, will be designed later in Section IV-C. Next, we provide the design of the vector $\mathbf{v}(\mathbf{x},\mathbf{a})\in\mathbb{R}^{3}$.
@@
$\mathbf{v}(\mathbf{x},\mathbf{a})=\big{[}\eta(\mathbf{x})\mathbf{I}+(1-|\eta(\mathbf{x})|)\mathbf{R}(\mathbf{a})\big{]}{{\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}}}$ (12)
where $\mathbf{I}\in\mathbb{R}^{3\times 3}$ is the identity matrix and $\mathbf{x}_{\pi}:=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ The operator $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ provides the closest point on the obstacle-occupied workspace from the location of the center of the robot $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\bar{r}_{s}-r_{s})$, according to (8), the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighborhood of these obstacles. On the other hand, there may be some locations in the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ for which the uniqueness of the closest point from robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 1, the design of the flow sets and jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
@@
where we used the fact that $(\mathbf{a}^{\times})^{2}=\mathbf{a}\mathbf{a}^{\top}-I$, with $\mathbf{a}^{\times}$ being the skew-symmetric matrix associated to the unit vector $\mathbf{a}=[a_{1}~{}a_{2}~{}a_{3}]^{\top}$, defined according to (1).
@@
Finally, the scalar function $\eta(\mathbf{x})\in[0,1]$ is given by
@@
where $0<\gamma_{a}<\gamma_{s}<\gamma.$ The scalar function $\eta$ is designed to ensure that the center of the robot remains inside the $\gamma-$neighborhood of the $r_{a}-$dilated obstacle-occupied workspace $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$ when it operates in the obstacle-avoidance mode in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$. This feature allows for the design of the jump set of the obstacle-avoidance mode, as discussed later in Section IV-B2, and ensures convergence to the target location, as stated later in Theorem 1. Next, we provide the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ used in (IV-A).
@@ IV-B Geometric construction of the flow and jump sets
Note that when the robot’s center at $\mathbf{x}$ belongs to the exit region, the line segment connecting its location to the target location $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})$ does not intersect with the interior of the nearest $r_{a}-$dilated obstacle. Hence, the robot should move straight towards the target location only if it is in the exit region. Next, we provide the geometric construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (IV-A).
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
According to (21) and (22), the robot operating in the obstacle-avoidance mode, can switch to the move-to-target mode only when its center belongs to the exit region $\mathcal{R}_{e}$ and is at least $\epsilon$ units closer to the target than the current hit point $\mathbf{h}.$
@@
Finally, the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (IV-A), are defined as
@@
where $\mathcal{F}_{0},\mathcal{J}_{0}$ defined in (20) for $m=0$ and in (24) for $m=1.$ Next, we provide the update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ used in (11o).
@@ IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m)$
The update law $\mathbf{L}(\xi)$, used in (11o), updates the value of the hit point $\mathbf{h}$, the unit vector $\mathbf{a}$, the mode indicator $m$ and the variable $s$ when the state $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ belongs to the jump set $\mathcal{J}$ defined in (25) and is given by
@@
$\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{q}\\ 1\\ s+1\end{bmatrix}\Bigg{|}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x})\right\},$ (27)
where for any $\mathbf{q}\in\mathbb{R}^{3},$ the set $\mathcal{P}^{\perp}(\mathbf{q})$, which is defined as
$\mathcal{P}^{\perp}(\mathbf{q}):=\{\mathbf{p}\in\mathbb{S}^{2}|\mathbf{p}^{\intercal}\mathbf{q}=0\},$ (28)
contains unit vectors that are perpendicular to the vector $\mathbf{q}.$
According to (27), when the robot switches from the move-to-target mode to the obstacle-avoidance mode, the coordinates of the hit point $\mathbf{h}$ get updated to the current value of $\mathbf{x}.$ Moreover, the unit vector $\mathbf{a}$ is updated such that it is perpendicular to the vector $\mathbf{x}$. This ensures that the target location $\mathbf{0}$ belongs to the hyperplane perpendicular to $\mathbf{a}$ passing through the hit point $\mathbf{h}$ which equals $\mathbf{x}$. Since the target location $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ and the hit point $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}$, the hyperplane intersects with both the landing region $\mathcal{R}_{l}$ and the exit region $\mathcal{R}_{e}$ associated with the nearest obstacle. This feature is then used to ensure that the robot, operating in the obstacle-avoidance mode, always enters in the move-to-target mode. This allows one to ensure global convergence to the target location, as stated later in Theorem 1.
@@
Since the parameter $\gamma\in(r_{s},\bar{r}_{s}-r_{s})$, where $\bar{r}_{s}$ is defined in (8), the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint, and the target location $\mathbf{0}\in\mathcal{W}_{r_{a}+\gamma}^{\circ}$. Furthermore, according to (23), the flow set of the obstacle-avoidance mode associated with the state $\mathbf{x},$ denoted as $\mathcal{F}_{1}^{\mathcal{W}}$, is contained within the region $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})).$ Hence, the proposed control law enables the robot to avoid one obstacle at a time.
@@
This concludes the design of the proposed hybrid feedback controller (IV-A). Next, we analyze the safety, stability and convergence properties of the proposed hybrid feedback controller.
@@ V Stability Analysis
The hybrid closed-loop system resulting from the hybrid feedback control law (IV-A) is given by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi)},{\xi\in\mathcal{F}},\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi)},{\xi\in\mathcal{J}},$ (30)
@@
    the flow map $\mathbf{F}$, defined in (30), is outer semicontinuous and locally bounded relative to $\mathcal{F}$, dom $\mathcal{F}\subset\text{dom }\mathbf{F},$ and $\mathbf{F}(\xi)$ is convex for every $\xi\in\mathcal{F},$

### v2

@@ II-A Notations
The sets of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We identify vectors using bold lowercase letters. The Euclidean norm of a vector $\mathbf{p}\in\mathbb{R}^{n}$ is denoted by $\lVert\mathbf{p}\rVert$, and an Euclidean ball of radius $r\geq 0$ centered at $\mathbf{p}$ is represented by $\mathcal{B}_{r}(\mathbf{p})=\{\mathbf{q}\in\mathbb{R}^{n}|\lVert\mathbf{q}-\mathbf{p}\rVert\leq r\}.$ The set of $n-$dimensional unit vectors is given by $\mathbb{S}^{n-1}=\{\mathbf{p}\in\mathbb{R}^{n}|\lVert\mathbf{p}\rVert=1\}.$ The identity matrix of order $n$ is denoted by $\mathbf{I}_{n}$. The three-dimensional Special Orthogonal group is defined by $SO(3):=\{\mathbf{R}\in\mathbb{R}^{3\times 3}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{3},\det(\mathbf{R})=1\}.$ For a given vector $\mathbf{x}:=[x_{1},x_{2},x_{3}]^{\top}\in\mathbb{R}^{3}$, we define $\mathbf{x}^{\times}$ as the skew-symmetric matrix, which is given by
@@ II-C Geometric subsets of $\mathbb{R}^{n}$
II-C1 Line
The line passing through two points $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{q}\in\mathbb{R}^{n}\setminus\{\mathbf{p}\}$ is given by
$\mathcal{L}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{x}=\lambda\mathbf{p}+(1-\lambda)\mathbf{q},\lambda\in\mathbb{R}\}.$ (4)
II-C2 Line segment
@@ IV-A Hybrid control design
$\underbrace{\begin{matrix}\mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}\;=\mathbf{0},\\ \;=0,\\ \;=0\\ \;=1,\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}}\quad\quad\underbrace{\begin{bmatrix}\mathbf{h^{+}}\\ \mathbf{a^{+}}\\ m^{+}\\ {s^{+}}\end{bmatrix}\begin{matrix}\;\in\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s),\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{J}}$ (12b)
where $\kappa_{s}>0$, $\kappa_{r}>0$, and $\xi=(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$. The variable $\mathbf{h}$ denotes the hit point, which is the location where the robot switches from the move-to-target mode to the obstacle-avoidance mode. The vector $\mathbf{a}\in\mathbb{S}^{2}$ is instrumental for the construction of the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$, used in (12a). The scalar variable $s\in\mathbb{R}_{\geq 0}$ allows the robot to switch from the obstacle-avoidance mode to the move-to-target mode when initially in the obstacle-avoidance mode. Details of this switching process are provided later in Section IV-B. The sets $\mathcal{F}$ and $\mathcal{J}$ are the flow and jump sets related to different modes of operation, respectively, whose constructions are provided in Section IV-B. The update law $\mathbf{L}$, which allows the robot to update the values of the variables $\mathbf{h}$, $\mathbf{a}$, $m$ and $s$ based on the current location of the robot with respect to the nearest obstacle and the target location, will be designed later in Section IV-C. Next, we provide the design of the vector $\mathbf{v}(\mathbf{x},\mathbf{a})\in\mathbb{R}^{3}$.
@@
$\mathbf{v}(\mathbf{x},\mathbf{a})=\big{[}\eta(\mathbf{x})\mathbf{I}_{3}+(1-|\eta(\mathbf{x})|)\mathbf{R}(\mathbf{a})\big{]}{{\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}}}$ (13)
where $\mathbf{I}_{3}\in\mathbb{R}^{3\times 3}$ is the identity matrix and $\mathbf{x}_{\pi}:=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ The operator $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ provides the closest point on the obstacle-occupied workspace from the location of the center of the robot $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\bar{r}_{s}-r_{s})$, according to (9), the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighborhood of these obstacles. On the other hand, there may be some locations in the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ for which the uniqueness of the closest point from robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 1, the design of the flow sets and jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
@@
where we used the fact that $(\mathbf{a}^{\times})^{2}=\mathbf{a}\mathbf{a}^{\top}-\mathbf{I}_{3}$, with $\mathbf{a}^{\times}$ being the skew-symmetric matrix associated to the unit vector $\mathbf{a}=[a_{1}~{}a_{2}~{}a_{3}]^{\top}$, defined according to (1).
@@
Finally, the scalar function $\eta(\mathbf{x})\in[-1,1]$ is given by
@@
where $0<\gamma_{a}<\gamma_{s}<\gamma.$ The scalar function $\eta$ is designed to ensure that the center of the robot remains inside the $\gamma-$neighborhood of the $r_{a}-$dilated obstacle-occupied workspace $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$ when it operates in the obstacle-avoidance mode in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$. This feature allows for the design of the jump set of the obstacle-avoidance mode, as discussed later in Section IV-B2, and ensures convergence to the target location, as stated later in Theorem 1. Next, we provide the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ used in (12).
@@ IV-B Geometric construction of the flow and jump sets
Note that when the robot’s center at $\mathbf{x}$ belongs to the exit region, the line segment connecting its location to the target location $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})$ does not intersect with the interior of the nearest $r_{a}-$dilated obstacle. Hence, the robot should move straight towards the target location only if it is in the exit region. Next, we provide the geometric construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (12).
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
According to (22) and (23), the robot operating in the obstacle-avoidance mode, can switch to the move-to-target mode when its center belongs to the exit region $\mathcal{R}_{e}$ and the target location $\mathbf{0}$ is at least $\epsilon$ units closer to $\mathbf{x}$ than the current hit point $\mathbf{h}$.
@@
Finally, the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (12), are defined as
@@
where $\mathcal{F}_{0},\mathcal{J}_{0}$ defined in (21) for $m=0$ and in (25) for $m=1.$ Next, we provide the update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ used in (12b).
@@ IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m)$
The update law $\mathbf{L}(\xi)$, used in (12b), updates the value of the hit point $\mathbf{h}$, the unit vector $\mathbf{a}$, the mode indicator $m$ and the variable $s$ when the state $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ belongs to the jump set $\mathcal{J}$ defined in (26) and is given by
@@
$\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\begin{bmatrix}\mathbf{x}\\ \mathbf{A}(\mathbf{x})\\ 1\\ s+1\end{bmatrix}.$ (28)
Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}:\mathbb{R}^{3}\rightrightarrows\mathbb{S}^{2}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\frac{\mathbf{x}^{\times}\mathbf{x}_{\pi}}{\lVert\mathbf{x}^{\times}\mathbf{x}_{\pi}\rVert},&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (29)
where for any $\mathbf{p}\in\mathbb{R}^{3}$, the set $\mathcal{P}^{\perp}(\mathbf{p})$, which is defined as
$\mathcal{P}^{\perp}(\mathbf{p}):=\{\mathbf{q}\in\mathbb{S}^{2}|\mathbf{q}^{\intercal}\mathbf{p}=0\},$ (30)
contains unit vectors that are perpendicular to the vector $\mathbf{p}.$
According to (28), when the robot switches from the move-to-target mode to the obstacle-avoidance mode, the coordinates of the hit point $\mathbf{h}$ get updated to the current value of $\mathbf{x}.$ Moreover, according to (29), the unit vector $\mathbf{a}$ is updated such that it is perpendicular to the vector $\mathbf{h}$ and the resulting hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$ intersects with the interior of the nearest $r_{a}-$dilated obstacle. Since the target location $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ and the hit point $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}$, the hyperplane intersects with both the landing region $\mathcal{R}_{l}$ and the exit region $\mathcal{R}_{e}$ associated with the nearest obstacle. This feature is then used to ensure that the robot, operating in the obstacle-avoidance mode, always enters in the move-to-target mode. This allows one to ensure global convergence to the target location, as stated later in Theorem 1.
@@
Since the parameter $\gamma\in(r_{s},\bar{r}_{s}-r_{s})$, where $\bar{r}_{s}$ is defined in (9), the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint, and the target location $\mathbf{0}\in\mathcal{W}_{r_{a}+\gamma}^{\circ}$. Furthermore, according to (24), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
@@
This concludes the design of the proposed hybrid feedback controller (12). Next, we analyze the safety, stability and convergence properties of the proposed hybrid feedback controller.
@@ V Stability Analysis
The hybrid closed-loop system resulting from the hybrid feedback control law (12) is given by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),{\xi\in\mathcal{F}}}\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),{\xi\in\mathcal{J}}},$ (32)
@@
    the flow map $\mathbf{F}$, defined in (32), is outer semicontinuous and locally bounded relative to $\mathcal{F}$, dom $\mathcal{F}\subset\text{dom }\mathbf{F},$ and $\mathbf{F}(\xi)$ is convex for every $\xi\in\mathcal{F}$.